# Scalable UAV Multi-Hop Networking via Multi-Agent Reinforcement Learning with Large Language Models

### v1

@@ V-A2 Training configuration
For the training process, a total of 25,000 episodes are conducted, with each episode consisting of 400 time slots. The PPO policy and critic networks are implemented as multi-layer perceptrons with five hidden layers, each employing the tanh activation function. And the learning rate is initially set to $3\times 10^{-4}$ and is progressively lowered to $1\times 10^{-4}$ toward the end of the training process. In equation (P1), the weighting factor $\kappa$ is defined as 0.025 when the data rate $r^{\text{UE}}$ is measured in Mbps. The coefficient $\beta_{1}$, which controls the weight of the knowledge distillation component, is initially set to 0.5 and is gradually reduced to 0.1 during the final phase of training. It is designed to help improve the effectiveness of MARL by gradually shifting the focus from distillation to the training objective. In MARL training, the UAVs are divided into four groups based solely on the distance relationship between each UAV and the set of BS nodes $\boldsymbol{\mathcal{G}}$, as defined by equation (19). The network parameters of each agent remain independent and are not influenced by the grouping strategy. Furthermore, a detailed discussion on the relationship between parameter sharing in groups and training performance can be found in Section V-D. Due to the substantial inference latency of the LLM, which adversely affects the training efficiency of the MARL agents, LLM inference is performed only at certain time slots in each episode to generate the necessary output. For subsequent time slots, the generated results are retained and used consistently in the computation of the knowledge distillation loss, $\mathcal{L}^{\text{KD}}$. Since the UE distribution typically experiences only minor changes over short time intervals, the retained outputs perform well during these periods. Additionally, given that the LLM agent is primarily tasked with guiding MARL agent training and managing an extensive exploration space, even suboptimal network decisions from the LLM are acceptable, ensuring that this intermittent LLM reasoning strategy does not lead to a significant negative impact on knowledge distillation. Additional critical parameter settings are summarized in Table I.
@@ V-A3 Performance metrics
The performance of the algorithm is assessed using two primary metrics: the average data rate $({1}/{M})\sum_{m=1}^{M}r^{\text{UE}}_{m}$ and the connected UE proportion $(1/{M})\sum_{m=1}^{M}c_{m}^{\text{UE}}$, i.e. the proportion of UEs that are successfully connected to the core network relative to the total number of UEs. To evaluate the robustness of each model in the networking task, the number of available UAVs is introduced as another metric. This metric quantifies the number of UAVs that successfully establish and maintain connectivity with the BSs and core network through the multi-hop UAV network. A higher value indicates fewer disconnections, reflecting a more resilient and stable network. Formally, it is measured as $(1/U)\sum_{u\in\boldsymbol{\mathcal{U}}}c_{u}^{\text{UAV}}$, where $c_{u}^{\text{UAV}}$ denotes the connectivity status of UAV $u$, defined in equation (11). This metric serves as a key indicator of the stability and effectiveness of each algorithm.
@@ V-B Comparative Experiments
In this section, we evaluate the performance of the proposed approach MRLMN by comparing three different methods:
@@
  • GA2C: A RL framework [46] that integrates graph neural network (GNN) to guide action training within an advantage actor-critic architecture. By leveraging hidden representations from network feature correlations, this method captures intricate environmental relationships, thereby enhancing the efficiency of policy training.
@@
Figure 4 shows the training curves over 10 million steps, depicting the performance trends of different methods in terms of the objective function defined in equation (P1). The experiments are set in a 3.5 km $\times$ 3.5 km environment with 18 UAVs, evaluating the proposed framework under large-scale networking conditions. The results show that the proposed MRLMN method significantly outperforms the baseline methods throughout the training process. MRLMN exhibits a rapid increase in training performance, stabilizing above 0.8, while the competing methods plateau at much lower values. Specifically, MAPPO and GA2C stabilize between 0.4 and 0.6, while GVis performs slightly better but remains well below MRLMN. The curves indicate that the proposed algorithm, by integrating multiple modules, achieves superior networking performance. Specifically, the knowledge distillation module enables reinforcement learning agents to rapidly acquire generalized decision-making capabilities, accelerating the overall training process. In parallel, the reward decomposition and behavioral constraints allow for fine-tuning of agent behaviors, ensuring a better adaptation to the unique challenges of UAV multi-hop networking.
Furthermore, the comparison is conducted across several metrics under varying network conditions, including changes in geographical area size and the number of deployed UAVs, providing a comprehensive view of scalability and network robustness.
@@ V-B1 Impact of environment size
Figure 5.(a), (c), (e) examine the influence of expanding the square environment area from 6.76 km2 to 14.44 km2, with the number of UAVs fixed at 18. The shaded regions indicate the standard deviation of each method, providing insight into algorithmic stability. In Figure 5.(a), the number of connected UEs for all approaches declines as the area expands, reflecting the increased difficulty of maintaining reliable connections over larger regions populated with numerous UEs. Nevertheless, compared to GVis, GA2C, and MAPPO, MRLMN achieves an average UE coverage improvement of approximately 27% and maintains greater performance stability across the entire range of area sizes. A similar trend is observed in Figure 5.(c), which reports the average data rate in Mbps under the same conditions. MRLMN not only sustains better coverage but also delivers higher throughput across various environment sizes. For these two metrics, GVis and GA2C exhibit similar performance, both relying on GNN-based strategies for UAV coordination. However, MRLMN incorporates scenario-specific grouping and reward mechanisms, allowing for more adaptive decision-making. Additionally, its knowledge distillation method is designed for large-scale environments, further enhancing overall performance. In Figure 5.(e), all methods exhibit a reduction in the number of available UAVs, highlighting the increasing challenge of sustaining multi-hop connectivity over larger regions. Nonetheless, MRLMN maintains a higher number of available UAVs and lower variance compared to other methods, demonstrating its enhanced stability and robustness in large-scale networking scenarios. In smaller environments, MRLMN achieves nearly 100% availability of UAVs, while in larger environments, it outperforms the best alternative by approximately 17%. This advantage stems from MRLMN’s relay-oriented reward design and connectivity-preserving behavioral constraints, which enable more effective multi-hop link management.
@@ VI Conclusion
This paper introduces the MRLMN framework, which integrates MARL and LLMs to optimize UAV networking in disaster response scenarios. To address the scalability of the networking problem and enhance coordination among UAVs, the framework incorporates agent grouping and reward decomposition modules. Behavioral constraints based on the grouping mechanism are further applied to ensure robust and stable network formation. Additionally, a knowledge distillation approach enables the transfer of high-level decision-making capabilities from LLMs to MARL agents, accelerating training and improving exploration efficiency. Simulation results demonstrate substantial performance gains in large-scale dynamic environments, confirming the adaptability and effectiveness of the proposed framework across diverse configurations. Further investigation into practical constraints such as energy consumption and network load balancing, along with validation through real-world deployment experiments, would be valuable for enhancing the framework’s applicability in real deployments.

### v2

@@ V-A2 Training configuration
For the training process, a total of 25,000 episodes are conducted, with each episode consisting of 400 time slots. The PPO policy and critic networks are implemented as multi-layer perceptrons with five hidden layers, each employing the tanh activation function. And the learning rate is initially set to $3\times 10^{-4}$ and is progressively lowered to $1\times 10^{-4}$ toward the end of the training process. The LLM guidance is provided using the GPT-4o model [1], which generates high-level strategic deployment suggestions during MARL offline training. In the networking task, we assume identical transmitter and receiver antenna gains, with UAV, UE, and BS antenna gains set to 0, 0, and 5 dBi, respectively. In equation (P1), the weighting factor $\kappa$ is defined as 0.025 when the data rate $r^{\text{UE}}$ is measured in Mbps. The coefficient $\beta_{1}$, which controls the weight of the knowledge distillation component, is initially set to 0.5 and is gradually reduced to 0.1 during the final phase of training. It is designed to help improve the effectiveness of MARL by gradually shifting the focus from distillation to the training objective. In MARL training, the UAVs are divided into four groups based solely on the distance relationship between each UAV and the set of BS nodes $\boldsymbol{\mathcal{G}}$, as defined by equation (19). The network parameters of each agent remain independent and are not influenced by the grouping strategy. Furthermore, a detailed discussion on the relationship between parameter sharing in groups and training performance can be found in Section V-D. Due to the substantial inference latency of the LLM, which adversely affects the training efficiency of the MARL agents, LLM inference is performed only at selected time slots within each episode, specifically once every $Q_{\text{LLM}}$ steps to obtain the required guidance output. For subsequent time slots, the validated LLM outputs are cached and subsequently reused in the computation of the knowledge distillation loss, $\mathcal{L}^{\text{KD}}$. Since the UE distribution typically experiences only minor changes over short time intervals, the retained outputs perform well during these periods. Additionally, given that the LLM agent is primarily tasked with guiding MARL agent training and managing an extensive exploration space, even suboptimal network decisions from the LLM are acceptable, ensuring that this intermittent LLM reasoning strategy does not lead to a significant negative impact on knowledge distillation. Additional critical parameter settings are summarized in Table I. To focus on evaluating MARL coordination and policy performance rather than detailed physical-layer optimization, the bandwidth allocation across BS-UAV, UAV-UAV, and UAV-UE links is assumed to be fixed, which serves to mitigate cross-link interference.
@@ V-A3 Performance metrics
The performance of the algorithm is assessed using two primary metrics: the quality of service, defined as the average data rate per UE, calculated as $({1}/{M})\sum_{m=1}^{M}r^{\text{UE}}_{m}$ and the connected UE proportion $(1/{M})\sum_{m=1}^{M}c_{m}^{\text{UE}}$, i.e. the proportion of UEs that are successfully connected to the core network relative to the total number of UEs. To evaluate the robustness of each model in the networking task, the number of available UAVs is introduced as another metric. This metric quantifies the number of UAVs that successfully establish and maintain connectivity with the BSs and core network through the multi-hop UAV network. A higher value indicates fewer disconnections, reflecting a more resilient and stable network. Formally, it is measured as $(1/U)\sum_{u\in\boldsymbol{\mathcal{U}}}c_{u}^{\text{UAV}}$, where $c_{u}^{\text{UAV}}$ denotes the connectivity status of UAV $u$, defined in equation (11). This metric serves as a key indicator of the stability and effectiveness of each algorithm.
@@ V-B Comparative Experiments
In this section, we present a comprehensive performance evaluation of the proposed MRLMN framework by benchmarking it against five representative baselines that include both independently trained agents and cooperative MARL approaches:
@@
  • GA2C: A RL framework [21] that integrates graph neural network (GNN) to guide action training within an Advantage Actor-Critic (A2C) architecture. By leveraging hidden representations from network feature correlations, this method captures intricate environmental relationships, thereby enhancing the efficiency of policy training.
@@
  • IA2C: An independent RL method based on the A2C architecture [23]. In this approach, each UAV agent learns its policy independently without explicitly considering the joint state or actions of other agents, serving as a baseline for non-cooperative learning.
  • MAA2C: A multi-agent variant of A2C that employs a centralized training and decentralized execution paradigm similar to MAPPO but relies on the A2C optimization objective.
Figure 4 shows the training curves over 10 million steps, depicting the performance trends of different methods in terms of the objective function defined in equation (P1). The experiments are set in a 3.5 km $\times$ 3.5 km environment with 18 UAVs, evaluating the proposed framework under large-scale networking conditions. The results show that the proposed MRLMN method significantly outperforms the baseline methods throughout the training process. MRLMN exhibits a rapid increase in training performance, stabilizing above 0.8, while the competing methods plateau at much lower values. In detail, GVis performs relatively better among the baselines but still falls short of MRLMN. MAPPO and GA2C stabilize between 0.4 and 0.6, whereas the A2C-based methods converge more slowly and achieve lower overall performance, fluctuating around 0.4. The curves indicate that the proposed algorithm, by integrating multiple modules, achieves superior networking performance. Specifically, the knowledge distillation module enables reinforcement learning agents to rapidly acquire generalized decision-making capabilities, accelerating the overall training process. In parallel, the reward decomposition and behavioral constraints allow for fine-tuning of agent behaviors, ensuring a better adaptation to the unique challenges of UAV multi-hop networking.
Furthermore, to provide a comprehensive view of scalability and network robustness, we conduct a detailed comparison against the key baselines (GVis, GA2C, and MAPPO) across several metrics under varying network conditions, including changes in geographical area size and the number of deployed UAVs.
@@ V-B1 Impact of environment size
Figure 5.(a), (c), (e) examine the influence of expanding the square environment area from 6.76 km2 to 14.44 km2, with the number of UAVs fixed at 18. The shaded regions indicate the standard deviation of each method, providing insight into algorithmic stability. In Figure 5.(a), the number of connected UEs for all approaches declines as the area expands, reflecting the increased difficulty of maintaining reliable connections over larger regions populated with numerous UEs. Nevertheless, compared to GVis, GA2C, and MAPPO, MRLMN achieves an average UE coverage improvement of approximately 27% and maintains greater performance stability across the entire range of area sizes. A similar trend is observed in Figure 5.(c), which reports the average data rate in Mbps under the same conditions. MRLMN not only sustains better coverage but also delivers higher throughput across various environment sizes. For these two metrics, GVis and GA2C exhibit similar performance, both relying on GNN-based strategies for UAV coordination. However, MRLMN incorporates scenario-specific grouping and reward mechanisms, allowing for more adaptive decision-making. Additionally, its knowledge distillation method is designed for large-scale environments, further enhancing overall performance. Moreover, it is observed that the objectives of connectivity coverage and communication data rate are correlated. Adjustments in UAV networking that improve network connectivity also tend to increase communication rates, reflecting the overall performance gains achieved through coordinated multi-hop networking. In Figure 5.(e), all methods exhibit a reduction in the number of available UAVs, highlighting the increasing challenge of sustaining multi-hop connectivity over larger regions. Nonetheless, MRLMN maintains a higher number of available UAVs and lower variance compared to other methods, demonstrating its enhanced stability and robustness in large-scale networking scenarios. In smaller environments, MRLMN achieves nearly 100% availability of UAVs, while in larger environments, it outperforms the best alternative by approximately 17%. This advantage stems from MRLMN’s relay-oriented reward design and connectivity-preserving behavioral constraints, which enable more effective multi-hop link management.
@@ VI Conclusion
This paper introduces the MRLMN framework, which integrates MARL and LLMs to optimize UAV networking in disaster response scenarios. To address the scalability of the networking problem and enhance coordination among UAVs, the framework incorporates agent grouping and reward decomposition modules. Behavioral constraints based on the grouping mechanism are further applied to ensure robust and stable network formation. Additionally, a knowledge distillation approach enables the transfer of high-level decision-making capabilities from LLMs to MARL agents, accelerating training and improving exploration efficiency. Simulation results demonstrate substantial performance gains in large-scale dynamic environments, confirming the adaptability and effectiveness of the proposed framework across diverse configurations. Future research directions include incorporating practical constraints such as UAV energy consumption, network load balancing, and UAV replacement mechanisms to ensure continuous and reliable operations in large-scale deployments. Enhancing communication reliability through interference management and physical-layer optimization can further support the efficient implementation of multi-hop aerial networks. At the same time, deeper integration of LLMs with MARL, for example, by replacing conventional MARL policies with more expressive decision models, has the potential to improve scalability and adaptability in complex environments. Complementing these developments with validation in real-world deployments would provide valuable insights into the operational feasibility and robustness of the proposed framework.